Circular No. 8775

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 W1 (GIBBS) A. R. Gibbs reports his discovery of a comet on Catalina Sky Survey

CCD images obtained with the 0.68-m Schmidt telescope, the object showing a fairly distinct nuclear condensation (discovery observation tabulated below); four combined 60-s images on the same night in 2′′.2 seeing show a coma of diameter 8′′, a main tail 30′′ long in p.a. 280^{o}, and a possible secondary 22′′ tail in p.a. 315^{o}. S. Larson reports that images obtained on Nov. 17.5 UT at Mount Lemmon show the comet having a 20′′ coma and a broad, 100′′ tail in p.a. 285^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD observers have reported on the object’s cometary appearance, including P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain telescope; images on Nov. 18.19–18.25 UT show an elongated coma or possible tail up to 15′′ long in p.a. 270^{o}) and J. G. Ries (McDonald Observatory, 0.76-m reflector; R-band images on Nov. 18.5 show a diffuse area around the object, apparently elongated 10′′ slightly north of due west).

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 16.40896  8^{h}59^{m}27.^{s}60  +3^{o}11′01′′.0  18.6

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-W22.

T = 2006 May 11.583 TT  ω = 236.^{o}258
Ω = 157.133 }2000.0
q = 1.66452 AU  i = 17.855

COMET P/2006 U5 (CHRISTENSEN) Additional astrometry has shown that this comet (cf. IAUC 8768) has

a short orbital period, the improved elements below coming from MPEC 2006-W09:

T = 2007 Jan. 23.878 TT  ω = 100.^{o}212
e = 0.33759  Ω = 4.887 }2000.0
q = 2.31757 AU  i = 3.426
a = 3.49871 AU n^{o} = 0.150606 P = 6.54 years
2006 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green